Payments of $1200 are made semi-annually into an investment. The investment earns interest of 4.25% compounded...

  1. Payments of $1200 are made semi-annually into an investment. The investment earns interest of 4.25% compounded quarterly for 10 years.
  1. Find the balance of the investment at the end of the 10 year period
  1. How much interest is earned
  2. What is the effective rate of interest on the investment

(using financial calculator method (BAII) by showing inputs within the calculator)
Answer #1

APR(semi-annually) = 2[(1 + 0.0425/4)2 - 1] = 4.273%

Calculating Future Value in 10 years,

Using TVM Calculation,

FV = [PV = 0, PMT = 1,200, N = 20, I = 0.04273/2]

FV = $29,556.45

Interest Earned = (29,556.45 - 20(1,200))

Interest Earned = $5,556.45

EAR = (1 + 0.0425/4)4 -1

EAR = 4.318%